### v5

# Beyond Pixels: Enhancing LIME with Hierarchical Features and Segmentation Foundation Models

Abstract
LIME (Local Interpretable Model-agnostic Explanations) is a well-known XAI framework for unraveling decision-making processes in vision machine-learning models. The technique utilizes image segmentation methods to identify fixed regions for calculating feature importance scores as explanations. Therefore, poor segmentation can weaken the explanation and reduce the importance of segments, ultimately affecting the overall clarity of interpretation. To address these challenges, we introduce the DSEG-LIME (Data-Driven Segmentation LIME) framework, featuring: i) a data-driven segmentation for human-recognized feature generation by foundation model integration, and ii) a user-steered granularity in the hierarchical segmentation procedure through composition. Our findings demonstrate that DSEG outperforms on several XAI metrics on pre-trained ImageNet models and improves the alignment of explanations with human-recognized concepts.
[…]
3235
1 Introduction
Why should we trust you? The integration of AI-powered services into everyday scenarios, with or without the need for specific domain knowledge, is becoming increasingly common. Consider autonomous driving or surveillance, where accurate visual recognition of images or video streams is critical. In such high-stakes applications, precision and alignment with expert judgment and safety standards are paramount. To ensure reliability and safety, stakeholders frequently evaluate AI performance after deployment. For example, one might assess whether an AI correctly identifies pedestrians or hazardous objects in real-time footage, accurately detecting potential threats to avoid accidents or breaches in security. The derived question - "Why should we trust the model?" - is a direct reference to Local Interpretable Model-agnostic Explanations (LIME) [49]. LIME seeks to demystify AI decision-making by identifying key features that influence the output of a model, underlying the importance of the Explainable AI (XAI) research domain, particularly when deploying opaque models in real-world scenarios [5, 20, 32].
Figure 1: DSEG vs. other segmentation techniques within LIME. LIME-generated explanations [49] for EfficientNetB4 [53] using various segmentation methods: DSEG (ours) with SAM [28], Quickshift [23], SLIC [1], Felzenszwalb [16], and Watershed [41]. The top predictions are ‘Acoustic Guitar’ ($p$ = 0.31), ‘Golden Retriever’ ($p$ = 0.24), and ‘Electric Guitar’ ($p$ = 0.07).
Ambiguous explanations. LIME relies on conventional superpixel segmentation (e.g., graph‑ or clustering‑based methods [55]) to isolate image regions for classification explanations. However, these regions often lack semantic coherence or compositional structure, such as distinguishing car doors, wheels, or headlights, which undermines both clarity and relevance. Because LIME defaults to these arbitrary segmentation methods [49], highlighted areas often conflict with human intuition [27, 38]. Moreover, segmentation granularity has a dramatic impact on explanation quality [50]: excessive segments introduce instability, causing LIME to produce mutually contradictory explanations for the same image and eroding trust in both the explanation mechanism and the model itself [4, 20, 54, 60, 61].
This work. We introduce DSEG-LIME (Data-driven Segmentation LIME), an enhanced LIME framework that replaces traditional superpixel algorithms with flexible foundation-model segmenters. Users can plug in general-purpose models such as SAM (Segment Anything) [28] or domain-adapted variants through SAM adapters [11], producing semantically coherent regions that mirror human concepts. Building on these powerful segmenters, we propose a hierarchical segmentation scheme that encodes a compositional object structure. This lets users dial in the granularity of explanations, viewing, for example, a car as a single unit or decomposed into doors, wheels, or headlights. By enabling concept-level analysis at multiple scales, DSEG-LIME delivers clearer, more relevant feature highlights and allows independent evaluation of each subconcept. As shown in Figure˜1, when applied to an image of a dog playing guitar, DSEG-LIME distinctly pinpoints meaningful parts (e.g., the dog’s snout, the guitar’s body), demonstrating superior alignment with human intuition compared to existing LIME variants.
Key contributions. We advance explainable image analysis by:
  • Introduce DSEG‑LIME, the first adaptation of LIME to harness foundation segmentation models, enabling semantically coherent region proposals.
  • Develop a hierarchical segmentation framework embedding a compositional object structure, empowering users with fine‑grained control over explanation granularity—from whole objects to individual parts (e.g., doors, wheels, headlights).
  • Evaluate DSEG‑LIME against conventional segmentation and LIME variants on multiple pretrained classifiers, employing quantitative fidelity metrics [40] and a qualitative user study to assess both explanatory accuracy and human interpretability, while highlighting gaps between human intuition and model reasoning [38, 19].
2 Related work
Region-based perturbation XAI techniques. LIME is among several techniques designed to explain black box models through image perturbation. Fong and Vedaldi [18] introduced a meta-predictor framework that identifies critical regions via saliency maps. Subsequently, Fong et al. [17] developed the concept of extremal perturbations to address previous methods’ limitations. Additionally, Kapishnikov et al. [24] advanced an integrated-gradient, region-based attribution approach for more precise model explanations. More recently, Escudero-Viñolo et al. [15] highlight the constraints of perturbation-based explanations, advocating for the integration of semantic segmentation to enhance image interpretation.
Instability of LIME. The XAI community widely recognizes the instability in LIME’s explanations, which stems from LIME’s design [4, 54, 60, 61]. Alvarez-Melis and Jaakkola [4] handle this issue by showing the instability of various XAI techniques when slightly modifying the instance to be explained. A direct improvement is Stabilized-LIME (SLIME) proposed by Zhou et al. [61] based on the central limit theorem to approximate the number of perturbations needed in the data sampling approach to guarantee improved explanation stability. Zhao et al. [60] improve stability by exploiting prior knowledge and using Bayesian reasoning - BayLIME. GLIME [54] addresses this issue by employing an improved local and unbiased data sampling strategy, resulting in explanations with higher fidelity - similar to the work by Rashid et al. [47]. Recent advancements include Stabilized LIME for Consistent Explanations (SLICE) [7], which improves LIME through a novel feature selection mechanism that removes spurious superpixels and introduces an adaptive perturbation approach for generating neighborhood samples. Another hierarchical-based variation, DLIME [57], utilizes agglomerative hierarchical clustering to organize training data, focusing on tabular datasets. In contrast, DSEG-LIME extends this concept to images by leveraging the hierarchical structure of image segments.
Segmentation influence on explanation. The segmentation algorithm utilized to sample data around the instance $\mathbf{x}$ strongly influences its explanation. It directly affects the stability of LIME itself, as suggested by Ng et al. [42]. This behavior is in line with the investigation by Schallner et al. [50] that examines the influence of different segmentation techniques in the medical domain, showing that the quality of the explanation depends on the underlying feature generation process. Bluecher et al. [6] explore how occlusion and sampling strategies affect model explanations when integrated with segmentation techniques for XAI, including LRP (Layer-Wise Relevance Propagation) [39] and SHAP [34]. Their study highlights how different strategies provide unique explanations while evaluating the SAM technique in image segmentation. Sun et al. [52] use SAM within the SHAP framework to provide conceptually driven explanations, which we discuss in Appendix B.4 [29]. This work reinforces the strategy of employing deeper models to elucidate simpler ones by seamlessly integrating these approaches into widely adopted XAI frameworks.
Segmentation hierarchy. Hierarchical Segment Grouping (HSG) [25] performs unsupervised hierarchical semantic segmentation by leveraging multiview cosegmentation and clustering transformers to enforce spatial and semantic consistency across images. The work of Li et al. [30] aims to simulate the way humans structure segments hierarchically and introduce a framework called Hierarchical Semantic Segmentation Networks (HSSN), which approaches segmentation through a pixel-wise multi-label classification task. HIPPIE (HIerarchical oPen-vocabulary, and unIvErsal segmentation), proposed by Wang et al. [56], extends hierarchical segmentation by merging text and image data multimodally. It processes inputs through decoders to extract and then fuse visual and text features into enhanced representations. While numerous hierarchical segmentation methods exist, their specialized, end‑to‑end pipelines cannot accommodate external foundation models—making them incompatible with model‑agnostic explainers like LIME. To bridge this gap, we introduce a lightweight, compositional hierarchy within LIME that supports plug‑and‑play integration of any foundation or domain‑specific segmenter.
3 Foundations of LIME
LIME is a prominent XAI framework to explain a neural network $f$ in a model-agnostic and instance-specific (local) manner. It applies to various modalities, such as images or text [49]. In the following, we will briefly review LIME’s algorithm for treating images, also visualized in Figure˜2 next to our adaptation with DSEG.
Figure 2: DSEG Pipeline in LIME. LIME image analysis pipeline with DSEG integrated into feature generation. Dashed lines indicate the option to apply DSEG. The frozen segmentation foundation model can be freely chosen by the user.
Notation. We consider the scenario in which we deal with image data. Let $\mathbf{x}\in\mathcal{X}$ represent an image within a set of images, and let $\mathbf{y}\in\mathcal{Y}$ denote its corresponding label in the output space with logits $\mathcal{Y}\subseteq\mathbb{R}$ indicating the labels in $\mathcal{Y}$. We denote the neural network we want to explain by $f:\mathcal{X}\rightarrow\mathcal{Y}$. This network functions by accepting an input $\mathbf{x}$ and producing an output in $\mathcal{Y}$, which signifies the probability $p$ of the instance being classified into a specific class.
Feature generation. The technique involves training a local, interpretable surrogate model $g\in G$, where $G$ is a class of interpretable models, such as linear models or decision trees, which approximates $f$’s behavior around an instance $\mathbf{x}$ [49]. This instance needs to be transformed into a set of features that can be used by $g$ to compute the importance score of its features. In the domain of imagery data, segmentation algorithms segment $\mathbf{x}$ into a set of superpixels $s_{0}...s_{d}\in\mathcal{S}^{D}$, done by conventional techniques [23]. We use these superpixels as the features for which we calculate their importance score. This step illustrates the motivation of Figure˜1, which underpins the quality of features affecting the explanatory quality of LIME.
Sample generation. For sample generation, the algorithm manipulates superpixels by toggling them randomly. Specifically, each superpixel $s_{i}$ is assigned a binary state, indicating this feature’s visibility in a perturbed sample $\mathbf{z}$. The presence (1) or absence (0) of these features is represented in a binary vector $\mathbf{z^{\prime}_{i}}$, where the $i$-th element corresponds to the state of the $i$-th superpixel in $\mathbf{z}$. When a feature $s_{i}$ is absent (i.e., $s_{i}=0$), its pixel values in $\mathbf{z}$ are altered. This alteration typically involves replacing the original pixel values with a non-information holding value, such as the mean pixel value of the image or a predefined value (e.g., black pixels) [49, 54].
Feature attribution. LIME employs a proximity measure, denoted as $\pi_{\mathbf{x}}$, to assess the closeness between the predicted outputs $f(\mathbf{z})$ and $f(\mathbf{x})$, which is fundamental in assigning weights to the samples. In the standard implementation of LIME, the kernel $\pi_{\mathbf{x}}(\mathbf{z})$ is as $\pi_{\mathbf{x}}(\mathbf{z^{\prime}})=\exp\left(-\frac{D(\mathbf{x^{\prime}},\mathbf{z^{\prime}})^{2}}{\sigma^{2}}\right)$, where $\mathbf{x^{\prime}}$ is a binary vector, all states are set to 1, representing the original image $\mathbf{x}$. $D$ represents the $L2$ distance, given by $D(\mathbf{x^{\prime}},\mathbf{z^{\prime}})=\sqrt{\sum_{i=1}^{n}(\mathbf{x^{\prime}}_{i}-\mathbf{z^{\prime}}_{i})^{2}}$ and $\sigma$ being the width of the kernel. Subsequently, LIME trains a linear model, minimizing the loss function $\mathcal{L}$, which is defined as:
$\mathcal{L}(f,g,\pi_{\mathbf{x}})=\sum_{\mathbf{z},\mathbf{z^{\prime}}\in\mathcal{Z}}\pi_{\mathbf{x}}(\mathbf{z})\cdot(f(\mathbf{z})-g(\mathbf{z^{\prime}}))^{2}$ (1)
In this equation, $\mathbf{z}$, and $\mathbf{z^{\prime}}$ are sampled instances of the perturbed data set $\mathcal{Z}$, and $g$ is the interpretable model being learned [49, 54]. Interpretability is derived primarily from the coefficients of $g$. These coefficients quantify the influence of each feature on the model’s prediction, with each coefficient’s magnitude and direction (positive or negative) indicating the feature’s relative importance. DSEG breaks up this fixed structure and allows for a repeated calculation of features based on the concept hierarchy and user preferences.
4 DSEG-LIME
In this section, we will present DSEG-LIME’s two contributions: first, the substitution of traditional feature generation with a data-driven segmentation approach (Section˜4.1), and second, the establishment of a hierarchical structure that organizes segments in a compositional manner (Section˜4.2).
4.1 Data-driven segmentation integration
DSEG-LIME expands the LIME feature generation phase by incorporating data-driven segmentation models, outperforming conventional graph- or cluster-based segmentation techniques in creating recognizable image segments across various domains. For the remainder of the paper, we mainly use SAM [28] due to its remarkable capability to segment images in diverse areas. However, in the appendix, we show that it can also be applied to other segmentation models. Figure˜2 illustrates DSEG’s integration into the LIME framework of Section˜3. The input image is first passed to a segmentation foundation model, which outputs a set of segments. Importantly, we only use the generated segments for the explanation process and do not incorporate any additional information from the foundation model. By redefining the feature basis, DSEG alters the loss in Equation˜1 and the proximity metric for sampling, often producing a surrogate $g$ that more faithfully approximates the original model $f$ at instance $\mathbf{x}$. However, the effect of DSEG is comparable to other segmentation methods such as SLIC, since the surrogate model relies solely on the resulting segments and does not incorporate any additional information from the underlying segmentation foundation models.
4.2 Hierarchical segmentation
The segmentation capabilities of foundation models like SAM, influenced by its design and hyperparameters, allow fine and coarse segmentation of an image [28]. These models can segment a human-recognized concept at various levels, from the entirety of a car to its components, such as doors or windshields. This multitude of segments enables the composition of a concept into its sub-concepts, creating a hierarchical segmentation.
Our proposed framework, DSEG, operates in successive stages. To harness this hierarchy within LIME, we introduce a depth parameter $d$ that allows users to specify the desired level of segmentation granularity for more personalized explanations. At $d=1$, we compute the importance scores for all segments (coarse) of the top level. Segments with top $k$ scores, a user‐defined value, are then subdivided into their immediate children, and at $d=2$ we recompute importance scores on these finer segments. This refine–and–rescore cycle continues, incrementing $d$ at each iteration, until the user’s maximum depth is reached or no further subdivisions exist. By allowing the surrogate model to "zoom in" iteratively on the most salient human‐recognizable parts, DSEG produces explanations that are both semantically rich and tailored to the user’s needs. Next, we detail each step of the DSEG pipeline (2), illustrate its intermediate outputs (3), and present the complete pseudocode for clarity (Section˜A.1).
Figure 3: Visualized DSEG pipeline. Image (1) serves as the initial input, leading to its automatic segmentation depicted in (2). The hierarchical tree generated from this segmentation is illustrated in (3b), and (3a) showcases the mask composed of first-order nodes. Image (4) displays the finalized mask created after eliminating empty spaces, which is fed back into the sample generation of LIME. Image (5) represents the resultant explanation within the DSEG-LIME framework. The image shows an instance classified as an ’Airliner’ ($p$ = 0.86) by EfficientNetB4. Node 23 (blue node) indicates the segment that represents the superpixel of the airliner.
Automatic mask generation. Masks, also called segments or superpixels, represent distinct regions of an image. In the following, we denote the segmentation foundation model, such as SAM, by $\zeta$. Depending on the foundation model employed, $\zeta$ can be prompted using various methods, including points, area markings, text inputs, or automatically segmenting all visible elements in an image. For the main experiments of DSEG, we utilize the last prompt, automated mask generation, since we want to segment the whole image for feature generation without human intervention. We express the process as follows:
$M_{\text{auto}}=\zeta(\mathbf{x},G_{\text{prompt}})\text{, with }\mathcal{S^{D}}=M_{\text{auto}},$ (2)
where $\mathbf{x}$ denotes the input image, and $G_{\text{prompt}}$ specifies a general prompt configuration designed to enable automated segmentation. The output, $M_{\text{auto}}$, represents the automatically generated mask, as shown in Figure˜3 (2). For this work, we used SAM with a grid overlay, parameterized by the number of points per side, to facilitate the automated segmentation process.
Small cluster removal. The underlying foundation model generates segments of varying sizes. We define a threshold $\theta$ such that segments with pixel-size below $\theta$ are excluded:
$\mathcal{S^{\prime}}=\{s_{i}\in\mathcal{S^{D}}\,|\,\text{size}(s_{i})\geq\theta\}.$ (3)
In this study, we set $\theta=500$ to reduce the feature set (see Section˜C.1). The remaining superpixels in $\mathcal{S^{\prime}}$ are considered for feature attribution. We incorporate this feature into DSEG to enable user-driven segment exclusion during post-processing, giving users control over the granularity within the segmentation hierarchy. This ensures that users can tailor the segmentation to their specific needs, thereby enhancing the method’s flexibility and adaptability.
Hierarchical ordering. To handle overlapping segments, we impose a tree hierarchical structure $\mathcal{T}=(\mathcal{V},\mathcal{E})$. In this structure, the overlap signifies that the foundation model has detected a sub-segment within a larger segment, representing the relationship between fine and coarse segments. The final output of DSEG, utilized for feature calculation, excludes overlapping segments. The nodes $v\in\mathcal{V}$ denote segments in $\mathcal{S^{\prime}}$, and the edges $(u,v)\in\mathcal{E}$ encode the hierarchical relationship between segments. This hierarchical ordering process $H(\mathcal{S^{\prime}})$ is a composition of the relative overlap of the segments, defined as:
$H(\mathcal{S^{\prime}})=\text{BuildHierarchy}(\mathcal{S^{\prime}},\text{OverlapMetric}),$ (4)
where OverlapMetric quantifies the extent of overlap between two segments $s_{1},s_{2}\in\mathcal{S^{\prime}}$ defined by
$\text{OverlapMetric}(s_{1},s_{2})=\frac{|s_{1}\cap s_{2}|}{|s_{2}|}.$ (5)
The hierarchy prioritizes parent segments (e.g., person) over child segments (e.g., clothing), as depicted (3a) in Figure˜3. Each node represents one superpixel with its unique identifier. Section˜A.2 presents the BuildHierarchy method in greater detail. The depth $d$ of the hierarchy determines the granularity of the explanation, as defined by the user. A new set $\mathcal{S}^{\prime}_{d}$, with $d=1$, includes all nodes below the root. For $d>1$, DSEG does not start from the beginning. Instead, it uses the segmentation hierarchy and segments $\mathcal{S^{\prime}}$ from the first iteration. It adds the child nodes of the top-$k$ most significant parent nodes in $\mathcal{S}’_{d}$ at depth d1d-1, identified during feature attribution. These are the only nodes used to train the surrogate model. We visualize this selection in the tree shown in Figure˜3 (3b), where all nodes with depth one, including the children of node 23, are considered in the second iteration. For the scope of this paper, we concentrate on the first-order hierarchy ($d=1$) but provide additional explanations with $d=2$ in Section˜B.8.
Empty space removal. In hierarchical segmentation, some regions occasionally remain unsegmented. We refer to these areas as $R_{\text{unseg}}$. To address this, we employ the nearest neighbor algorithm, which assigns each unsegmented region in $R_{\text{unseg}}$ to the closest segment within the set $\mathcal{S}^{\prime}_{d}$:
$\mathcal{S}_{d}=\text{NearestNeighbor}(R_{\text{unseg}},\mathcal{S}^{\prime}_{d}).$ (6)
Although this modifies the distinctiveness of concepts, it enhances DSEG-LIME’s explanatory power. DSEG then utilizes the features $s_{0},\ldots,s_{d}\in\mathcal{S}_{d}$ for feature attribution within LIME. Figure˜3 (4) shows the corresponding mask along with the explanation of $d=1$ in step (5) for the ’airliner’ class. An ablation study of these steps is in Section˜C.1 and in Section˜C.7 we show exemplary feature attribution maps.
5 Evaluation
In the following section, we will outline our experimental setup (Section˜5.1) and introduce the evaluation framework designed to assess DSEG-LIME both quantitatively (Section˜5.2) and qualitatively (Section˜5.3), compared to other LIME methodologies utilizing various segmentation algorithms. Subsequently, we discuss the limitations of DSEG (Section˜5.4)
5.1 Experimental setup
Segmentation algorithms. Our experiment encompasses, along with SAM (vit_h), four conventional segmentation techniques: Simple Linear Iterative Clustering (SLIC) [1], Quickshift (QS) [23], Felzenszwalb (FS) [16] and Watershed (WS) [41]. We carefully calibrate the hyperparameters (e.g., kernel size, marker placement, number of segments, minimum segment size) to produce segment counts comparable to those generated by SAM. This calibration ensures that no technique is unfairly advantaged due to a specific segment count, for instance, scenarios where fewer but larger segments might yield better explanations than many smaller ones. In the supplementary material, we demonstrate the universal property of integrating other segmentation foundation models within DSEG by presenting additional experiments with DETR [8] and SAM 2 [48] in Section˜B.6, Section˜B.7.
Models to explain. The models investigated in this paper rely on pre-trained models, as our primary emphasis is on explainability. We chose EfficientNetB4 and EfficientNetB3 [53] as the ones treated in this paper, where we explain EfficientNetB4 and use EfficientNetB3 for a contrastivity check [40] (Section˜5.2.1). To verify that our approach works on arbitrary pre-trained models, we also evaluated it using ResNet-101 [22, 36] (Section˜B.1), VisionTransformer (ViT-384) [14] (Section˜B.2), and ConvNext (Tiny-224) (Section˜B.3)[33]. Furthermore, we demonstrate the applicability of our approach on a zero-shot learning example of CLIP [44] using a new dataset with other classes (Section˜C.5).
Dataset. We use images from the ImageNet classes [13], on which the covered models were trained [14, 22, 53]. Our final dataset consists of 100 carefully selected instances (Section˜D.1), specifically chosen to comprehensively evaluate the techniques quantitatively. We aligned with this number to maintain consistency with the sample size used in previous studies [52, 54]. However, we want to emphasize that the selection of images is not biased toward any model. We also test the approach for another dataset in Section˜C.5.
Table 1: Quantitative summary - classes. The table presents four quantitative areas and their metrics, comparing five segmentation techniques applied to EfficientNetB4: DSEG with SAM and comparative methods SLIC, Quickshift (QS), Felzenszwalb’s (FS), and Watershed (WS). We test each with four LIME framework variations: LIME (L), SLIME (S), GLIME (G), and BayLIME (B). The experimental setup and metrics are detailed in Section˜5.1 and Section˜5.2.1. The table shows class metrics, each with a max score of 100. Arrows indicate performance improvement, highlighting the best scores for each metric bold.
Domain Metric DSEG SLIC QS FS WS
L S G B L S G B L S G B L S G B L S G B
Correctness Random Model $\uparrow$ 74 74 74 74 68 67 68 68 70 70 70 70 69 70 70 70 67 66 67 66
Random Expl. $\uparrow$ 86 90 93 88 81 79 80 84 76 82 75 81 87 83 85 81 79 81 86 85
Single Deletion $\uparrow$ 63 62 64 63 36 34 33 34 19 20 17 18 29 30 27 30 24 27 26 24
Output Completeness Preservation $\uparrow$ 77 74 73 73 74 74 75 74 68 65 67 64 71 74 74 72 79 79 77 78
Deletion $\uparrow$ 72 74 74 74 39 39 39 40 33 34 35 34 43 43 43 43 44 44 44 44
Consistency Noise Preservation $\uparrow$ 76 74 77 77 72 72 72 72 58 57 60 57 62 64 63 63 71 70 71 71
Noise Deletion $\uparrow$ 68 66 66 66 40 41 40 40 32 34 33 34 39 41 39 41 48 49 48 49
Contrastivity Preservation $\uparrow$ 64 63 64 65 54 54 55 55 49 46 46 45 49 52 53 52 54 54 54 53
Deletion $\uparrow$ 69 70 71 71 49 50 48 50 39 39 40 41 42 42 42 42 47 47 47 47
Hyperparameters and hardware setup. The experiments were conducted on an Nvidia RTX A6000 GPU. We compare standard LIME, SLIME [61], GLIME [54], and BayLIME [60], all integrated with DSEG, using 256 samples per instance, a batch size of ten and mean superpixel value for perturbation. For each explanation, up to three features are selected based on their significance, identified by values that exceed the average by more than 1.5 times the standard deviation. In BayLIME, we use the ‘non-info-prior’ setting to avoid introducing priors benefiting one method, even though the original papers show that this can have a diminishing effect. For SAM, we configure it to use 32 points per side, and conventional segmentation techniques are adjusted to achieve a similar segment count, as previously mentioned. In SLIC, we modify the number of segments and compactness; in Quickshift, the kernel size and maximum distance; in Felzenszwalb, the scale of the minimum size parameter; and in Watershed, the number of markers and compactness. Other hyperparameters remain at default settings to ensure a balanced evaluation across methods. Additional comparisons with SLICE [7] and EAC [52] are in the supplementary material.
5.2 Quantitative evaluation
We adapt the framework by Nauta et al. [40] to quantitatively assess XAI outcomes in this study, covering three domains: content, presentation, and user experience. In the content domain, we evaluate correctness, output completeness, consistency, and contrastivity. Presentation domain metrics like compactness and confidence are assessed under content for simplicity. We will briefly describe each metric individually to interpret the results correctly. The user domain, detailed in Section˜5.3, includes a user study that compares our approach with other segmentation techniques in LIME. We use quantitative and qualitative assessments to avoid over-emphasizing technical precision or intuitive clarity [38].
5.2.1 Quantitative metrics definition
Correctness involves two randomization checks. The model randomization parameter check (Random Model) [2] tests if changing the random model parameters leads to different explanations. The explanation randomization check (Random Expl.) [35] examines if random output variations in the predictive model yield various explanations. For both metrics, in Table˜1 we count the instances where explanations result in different predictions when reintroduced into the model under analysis. The domain also utilizes two deletion techniques: single deletion [3] and incremental deletion [21, 23]. Single deletion serves as an alternative metric to assess the completeness of the explanation, replacing less relevant superpixels with a specific background to evaluate their impact on the model predictions [45]. After these adjustments, we note instances where the model maintains the correct image classification. Incremental deletion (Incr. Deletion) entails progressively eliminating features from most to least significant based on their explanatory importance. We observe the model’s output variations, quantifying the impact by measuring the area under the curve (AUC) of the model’s confidence, as parts of the explanation are excluded. This continues until a classification change is observed (not ground truth class), and the mean AUC score for this metric is documented in Table˜2.
Output completeness measures whether an explanation covers the crucial area for accurate classification. It includes a preservation check (Preservation) [21] to assess whether the explanation alone upholds the original decision, and a deletion check (Deletion) [59] to evaluate the effect of excluding the explanation on the prediction outcome [45]. This approach assesses both the completeness of the explanation and its impact on the classification. The results are checked to ensure that the consistency of the classification is maintained. Compactness is also considered, highlighting that the explanation should be concise and cover all the areas necessary for prediction [9], reported by the mean value.
Consistency assesses explanation robustness to minor input alterations, like Gaussian noise addition, by comparing pre-and post-perturbation explanations for stability against slight changes (Noise Stability) [58], using both preservation and deletion checks. For consistency of the feature importance score, we generate explanations for the same instance 16 times (Rep. Stability), calculate the standard deviation $\sigma_{i}$ for each coefficient $i$, and then average all $\sigma_{i}$ values. This yields $\bar{\sigma}$, the average standard deviation of coefficients, and is reported as the mean score. Furthermore, we reference [7] and compute a Gini coefficient to assess the inequality of the coefficients of the superpixels. We also evaluate DSEG directly within SLICE in Section˜B.5.
Contrastivity integrates several previously discussed metrics, aiming for target-discriminative explanations. This means that an explanation $e_{\textbf{x}}$ for an instance x from a primary model $f_{1}$ (EfficientNetB4) should allow a secondary model $f_{2}$ (EfficientNetB3) to mimic the output of $f_{1}$ as $f_{1}(\textbf{x})\approx f_{2}(e_{\textbf{x}})$ [51]. The approach checks the explanation’s utility and transferability across models, using EfficientNetB3 for preservation and deletion tests to assess consistency.
Table 2: Quantitative summary - numbers. The table summarizes metrics from Section˜5.2.1, focusing on those quantified by rational numbers like incremental deletion, compactness, representational stability, and average computation time across the examples, detailed in Section˜5.1. Once more, the arrows denote the direction of improvement performance.
Metric DSEG SLIC QS FS WS
L S G B L S G B L S G B L S G B L S G B
Gini $\uparrow$ 0.53 0.52 0.54 0.53 0.49 0.49 0.50 0.50 0.41 0.42 0.41 0.41 0.45 0.44 0.45 0.45 0.46 0.47 0.48 0.47
Incr. Deletion $\downarrow$ 1.12 0.44 0.45 0.50 0.81 0.82 0.81 0.82 1.57 1.58 1.60 1.61 1.49 1.48 1.54 1.52 0.79 0.80 0.79 0.78
Compactness $\downarrow$ 0.17 0.18 0.18 0.18 0.15 0.15 0.15 0.15 0.12 0.12 0.12 0.12 0.12 0.12 0.13 0.13 0.12 0.12 0.12 0.13
Rep. Stability $\downarrow$ .010 .010 .010 .010 .011 .011 .012 .012 .011 .011 .012 .011 .012 .011 .012 .011 .012 .012 .012 .012
Time $\downarrow$ 28.5 31.1 31.9 31.7 16.1 16.2 16.8 16.8 28.6 28.6 28.9 28.6 15.9 16.0 17.3 16.7 15.7 15.6 17.0 16.5
5.2.2 Quantitative evaluation results
Table˜1 presents the outcomes of all metrics associated with class-discriminative outputs. The numbers in bold signify the top results, with an optimal score of 100. We compare LIME (L) [49] with the LIME techniques discussed in Section˜2, SLIME (S) [61], GLIME (G) [54], and BayLIME (B) [60] in combination with DSEG and the segmentation techniques from Section˜5.1. The randomization checks in the correctness category confirm that the segmentation algorithm bias does not inherently affect any model. This is supported by the observation that most methods correctly misclassify when noise is introduced or the model’s weights or predictions are shuffled. In contrast, DSEG excels in other metrics, surpassing alternative methods regardless of the LIME technique applied. In the output completeness domain, DSEG’s explanations more effectively capture the critical areas necessary for the model to accurately classify an instance, whether by isolating or excluding the explanation. This efficacy is supported by the single deletion metric, akin to the preservation check but with a perturbed background. Moreover, noise does not compromise the consistency of DSEG’s explanations. The contrastivity metric demonstrates DSEG’s effectiveness in creating explanations that allow another AI model to produce similar outputs in over half of the cases and outperform alternative segmentation approaches. Overall, the impact of the LIME feature attribution calculation remains relatively consistent, as we possess an average of 21.07 segments in the dataset under evaluation.
Table˜2 further illustrates DSEG’s effectiveness in identifying key regions for model output, especially in scenarios of incremental deletion where GLIME outperforms. Bolded values represent the best performance in their according category. Although compactness metrics show nearly uniform segment sizes across the techniques, smaller segments do not translate to better performance in other areas. Repeated experimentation suggests that stability is less influenced by the LIME variant (with the chosen hyperparameters) and more by the segmentation approach. In addition, the Gini value highlights that the feature values are more distinct compared to others. Further experiments show that DSEG outperforms the other techniques regarding stability as the number of features increases. This advantage arises from the tendency of data-driven approaches to represent known objects uniformly as a single superpixel (Section˜C.4). Thus, if a superpixel accurately reflects the instance that the model in question predicts, it can be accurately and effortlessly matched with one or a few superpixels - such accurate matching leads to a more precise and more reliable explanation. Conventional segmentation algorithms often divide the same area into multiple superpixels, creating unclear boundaries and confusing differentiation between objects. DSEG has an average runtime of approximately 30 seconds across different settings. Although the standard LIME approach using SLIC operates at around 16 seconds, the additional computational steps in DSEG (e.g., SAM integration and hierarchical ordering) introduce only a modest overhead. Importantly, when compared to the widely used Quickshift option, which typically takes about 29 seconds, DSEG’s runtime is directly comparable. Thus, the extra computational cost is minimal relative to the substantial benefits in explanation quality.
The supplementary experiments (see Appendix˜B) exhibit similarly consistent quantitative gains. We also extend our analysis by benchmarking against a SHAP‑based explainer [52], integrating DSEG into the SLICE framework [7], and testing different segmentation models.
5.3 Qualitative evaluation
User study. Following the methodology by Chromik and Schuessler [12], we conducted a user study (approved by the institute’s ethics council) to assess the interpretability of the explanations. This study involved 87 participants recruited via Amazon Mechanical Turk (MTurk) and included 20 random images in our dataset (Section˜D.1). These images were accompanied by explanations using DSEG and other segmentation techniques within the LIME framework (Section˜5.1).
Table 3: User study results. This table summarizes each segmentation approach’s average scores and top-rated counts of the user study results.
Metric DSEG SLIC QS FS WS
Avg. Score $\uparrow$ 4.16 3.01 1.99 3.25 2.59
Best Rated $\uparrow$ 1042 150 90 253 205
Participants rated the explanations on a scale from 1 (least effective) to 5 (most effective) based on their intuitive understanding and the predicted class. Table˜3 summarizes the average scores, the cumulative number of top-rated explanations per instance, and the statistical significance of user study results for each segmentation approach. DSEG is most frequently rated as the best and consistently ranks high, even when it is not the leading explanation. Paired t-tests indicate that DSEG is statistically significantly superior (additional results in Section˜D.2).
Exemplary Explanations. Figure˜4 displays five examples from our evaluation, with the top images showing DSEG’s explanations at a hierarchy depth of one and the bottom row at a depth of two. These explanations demonstrate that deeper hierarchies focus on smaller regions. However, the banana example illustrates a scenario where no further segmentation occurs if the concept, like a banana, lacks sub-components for feature generation, resulting in identical explanations at both depths. An example with a depth of three applied to a medical CT image is presented in Section˜C.2.
Figure 4: DSEG examples. Exemplary images from the evaluation dataset, illustrating DSEG explanations at $d=1$, and $d=2$.
These five images (with $d=1$) were also included in the user study. For the complete set of 100 evaluation images and their attribution maps, please refer to the supplementary material (Section˜C.7).
5.4 Limitations and future work
DSEG-LIME performs feature generation directly on images before inputting them into the model for explanation. For models like ResNet with smaller input sizes [22], the quantitative advantages are less evident (Section˜B.1). Experiments have shown that better results can be achieved with a lower SAM stability score threshold. Furthermore, substituting superpixels with a specific value in preservation and deletion evaluations can introduce an inductive bias [20]. To reduce this bias, using a generative model to synthesize replacement areas could offer a more neutral alteration. Additionally, future work should thoroughly evaluate feature attribution maps to ensure that methods assign significant attributions to the correct regions, as in Section˜C.7. This comprehensive assessment is essential to verify the interpretability and reliability of such methods. Lastly, our approach, like any other LIME-based method [49, 54, 60, 61], does not assume a perfect match between the explanation domains and the model’s actual domains since it simplifies the model by a local surrogate. Nonetheless, our quantitative analysis confirms that the approximations closely reflect the model’s behavior. Future work could focus on integrating foundation models directly into the system through a model-intrinsic approach, similar to [52].
No free lunch. While DSEG shows promising results across various domains, it is not universally applicable—its performance depends on the underlying segmentation model. If the model fails, for instance, due to a lack of domain-specific knowledge or the complexity of the feature generation task, the resulting concept extraction can be suboptimal (Section˜C.6). To address this limitation, future work could explore alternative segmentation backbones, such as HSSN [56] or HIPPIE [30], in place of SAM (or DETR). Moreover, recent extensions of SAM have introduced domain-adaptive modules that can improve segmentation in specialized contexts [11].
6 Conclusion
In this study, we introduced DSEG-LIME, an extension to the LIME framework, incorporating segmentation foundation models for feature generation with hierarchical feature calculation. This approach ensures that the generated features more accurately reflect human-recognizable concepts, enhancing the interpretability of explanations. Furthermore, we refined the process of feature attribution within LIME through an iterative method, establishing a segmentation hierarchy that contains the relationships between components and their subcomponents. In Section˜C.3, we show that our idea also helps explain a model’s wrong classifications. Through a comprehensive two-part evaluation, split into quantitative and qualitative analyses, DSEG emerged as the superior method, outperforming other LIME-based approaches in most evaluated metrics. The adoption of foundational models marks a significant step toward enhancing post-hoc and model-agnostic interpretability of deep learning models.
{ack}
This research was supported in part by the German Federal Ministry for Economic Affairs and Climate Action of Germany (BMWK), and in part by the German Federal Ministry for Research, Technology, and Space (BMFTR).
Appendix A Algorithms
A.1 DSEG-LIME algorithm
We present the pseudocode of our DSEG-LIME framework in Algorithm 1. To construct the hierarchical segmentation within our framework, we start by calculating the overlaps between all segments in $\mathcal{S}$. We build a hierarchical graph using this overlap information through the following process.
Algorithm 1 DSEG-LIME framework pseudocode
1:Input: $f$ (black-box model), $\zeta$ (segmentation function), $x$ (input instance), $g$ (interpretable model), $d$ (maximum depth), hphp (segmentation hyperparameters), $\theta$ (minimum segment size), $k$ (top segments to select)
2:1. Initial segmentation:
[…]
4:2. Small cluster removal:
[…]
6:3. Hierarchical ordering:
[…]
8:for $l\leftarrow 1$ to $d$ do
9:  4. Empty space removal:
10:  if $l=1$ then
11:   $\mathcal{S}_{l}\leftarrow\mathcal{H}[l]$
12:  else
13:   $\mathcal{S}_{l}\leftarrow\{s_{i}\in\mathcal{H}[l]\,|\,\text{parent}(s_{i})\in top\_ids\}$
14:  end if
15:  $\mathcal{S}_{l}\leftarrow\text{NearestNeighbor}(\mathcal{S}_{l})$
16:  $Z\leftarrow\text{Perturb}(x,\mathcal{S}_{l})$
17:  $w\leftarrow\text{Proximity}(Z,x)$
18:  $\text{preds}\leftarrow f(z)\text{ for all }z\in Z$
19:  $g\leftarrow\text{InitializeModel}(g)$
20:  $g\leftarrow\text{Fit}(g,Z,\text{preds},w)$
21:  $top\_ids\leftarrow\{\text{id}(s_{i})\,|\,s_{i}\in\mathcal{S}_{l},s_{i}\text{ is among top }k\text{ features in }g\}$
22:end for
23:Return $g$
First, we identify the top-level segments, which do not occur as subparts of any other segments. These segments serve as the highest-level nodes in the hierarchical graph. Starting from these top-level segments, we apply a top-down approach to identify child segments recursively. For each parent segment, we check for segments within it; these segments are designated as child nodes of the parent in the graph. Only the selected nodes are exclusively used to train the surrogate model.
This recursive process continues for each subsequent level, ensuring that every parent node encompasses its child nodes. The hierarchical graph thus formed represents the structural relationships between segments, where parent-child relationships indicate that child segments are complete parts of their respective parent segments. By constructing the hierarchy in this manner, we capture the nested structure of segments, which supports multi-level interpretability within the DSEG-LIME framework.
A.2 Hierarchical composition
The algorithm, designed for hierarchical composition, constructs a graph to represent the segments of a given instance, as shown in Algorithm˜2. A key feature of this algorithm is the additional hyperparameter $t$, which determines when one segment is considered a subpart of another based on their overlap. Initially, the graph includes loops and redundant edges. To ensure it accurately represents the hierarchical relationships, the algorithm applies a series of predefined rules to tune the graph, removing unnecessary connections and simplifying its structure for the final representation.
Algorithm 2 DSEG-LIME Hierarchical Compostion
1:Input: $\mathcal{S^{\prime}}$ (Segments), $t$ (Threshold)
2:1. Build overlap matrix:
3: Initialize overlap matrix $M$ of size $\|\mathcal{S^{\prime}}\|\times\|\mathcal{S^{\prime}}\|$ with zeros
4: Initialize graph $G$ with nodes corresponding to $\mathcal{S^{\prime}}$
5:for $i\leftarrow 1$ to $\|\mathcal{S^{\prime}}\|$ do
6:  for $j\leftarrow i$ to $\|\mathcal{S^{\prime}}\|$ do
7:   $M[i,j]\leftarrow\frac{|\mathcal{S^{\prime}}_{i}\cap\mathcal{S^{\prime}}_{j}|}{|\mathcal{S^{\prime}}_{j}|}$
8:   if $M[i,j]<t$ then
9:    $M[i,j]\leftarrow 0$
10:   else
11:    Add edge $(i,j)$ to $G$ with weight $M[i,j]$
12:   end if
13:  end for
14:end for
15:2. Remove loops in graph:
16:for each node $v\in G$ do
17:  if edge $(v,v)$ exists then
18:   Remove edge $(v,v)$
19:  end if
20:end for
21:3. Prune graph:
22:for each node $v\in G$ do
23:  Find all paths from $v$ to other nodes
24:  for each node $u$ reachable from $v$ via multiple paths do
25:   Identify the shortest path with the highest total weight
26:   If multiple paths have the same weight, select the longest one
27:   Remove all other paths from $v$ to $u$
28:  end for
29:end for
30:4. Create dictionary from graph:
31: Initialize dictionary $D$
32:for each node $v\in G$ do
33:  $D[v]\leftarrow\text{List of nodes connected to }v$
34:end for
35:Return $D$
Appendix B Supplementary model evaluations
B.1 ResNet
In Table˜4, we detail the quantitative results for the ResNet-101 model, comparing our evaluation with the criteria used for EfficientNetB4 under consistent hyperparameter settings. The review encompasses a comparative analysis with EfficientNetB3, emphasizing performance under contrastive conditions. The findings substantiate the results obtained from EfficientNet, demonstrating that the LIME techniques exhibit unpredictable behavior in the presence of model noise or prediction shuffling, despite varied segmentation strategies. This suggests an inherent randomness in the model explanations. Furthermore, the results indicate that all XAI methodologies displayed performance levels that were inferior to those of EfficientNetB4. However, it is significant to note that DSEG surpassed all other methods in terms of performance.
Table 4: Quantitative summary - classes ResNet-101. The table presents the metrics consistently with those discussed for EfficientNet.
Domain Metric DSEG SLIC QS FS WS
L S G B L S G B L S G B L S G B L S G B
Correctness Random Model $\uparrow$ 85 87 96 92 85 85 86 80 92 91 92 91 91 90 92 90 81 88 92 91
Random Expl. $\uparrow$ 95 93 97 94 94 92 94 91 94 94 92 95 96 94 97 97 91 95 98 99
Single Deletion $\uparrow$ 20 19 19 24 10 11 11 9 4 3 4 5 10 14 9 10 9 9 14 11
Output Completeness Preservation $\uparrow$ 48 39 38 39 42 45 40 37 26 27 28 34 37 35 36 36 29 38 39 38
Deletion $\uparrow$ 56 50 50 50 40 42 42 40 31 30 35 36 39 32 35 34 31 38 30 41
Consistency Noise Preservation $\uparrow$ 45 39 42 36 38 40 41 35 26 24 23 26 28 30 30 30 28 37 32 33
Noise Deletion $\uparrow$ 57 58 52 50 42 45 43 39 33 29 32 32 30 29 31 41 33 35 29 37
Contrastivity Preservation $\uparrow$ 45 39 39 43 36 35 36 35 31 37 30 40 42 43 43 39 42 41 44 41
Deletion $\uparrow$ 47 47 47 48 41 45 44 46 31 34 30 37 33 34 34 37 30 36 31 33
Table˜5 presents further findings of ResNet. SLIME with DSEG yields the lowest AUC for incremental deletion, whereas Quickshift and Felzenszwalb show the highest. WS produces the smallest superpixels for compactness, contrasting with DSEG’s larger ones. The stability analysis shows that all segmentations are almost at the same level, with QS being the best and GLIME the best-performing overall. Echoing EfficientNet’s review, segmentation defines runtime, with DSEG being the most time-consuming.
Table 5: Quantitative summary - numbers ResNet-101. The table presents the metrics consistently with those discussed for EfficientNet.
Metric DSEG SLIC QS FS WS
L S G B L S G B L S G B L S G B L S G B
Gini $\uparrow$ 0.43 0.44 0.45 0.44 0.41 0.42 0.42 0.41 0.38 0.39 0.38 0.40 0.41 0.40 0.41 0.39 0.41 0.41 0.42 0.40
Incr. Deletion $\downarrow$ 0.59 0.32 0.34 0.36 0.59 0.57 0.56 0.61 0.97 0.98 0.92 0.93 0.96 0.95 0.92 0.93 0.52 0.56 0.54 0.54
Compactness $\downarrow$ 0.22 0.23 0.22 0.22 0.17 0.17 0.17 0.17 0.14 0.13 0.13 0.13 0.15 0.15 0.15 0.15 0.13 0.13 0.12 0.13
Rep. Stability $\downarrow$ .021 .021 .018 .021 .019 .019 .016 .019 .018 .018 .015 .018 .018 .018 .016 .017 .019 .019 .016 .019
Time $\downarrow$ 8.0 8.2 8.1 8.4 2.5 2.5 2.4 2.4 11.4 11.5 11.5 11.4 2.8 2.7 2.8 2.9 2.9 2.9 3.2 22
B.2 VisionTransformer
Table˜6 provides the quantitative results for the VisionTransformer (ViT-384) model, employing settings identical to those used for EfficientNet and ResNet, with ViT processing input sizes of (384x384). The class-specific results within this table align closely with the performances recorded for the other models, further underscoring the effectiveness of DSEG. This consistency in DSEG performance is also evident in the data presented in Table˜7.
We performed all experiments for ResNet and ViT with the same hyperparameters defined for EfficientNetB4. We would like to explicitly point out that the quantitative results could be improved by defining more appropriate hyperparameters for both DSEG and conventional segmentation methods, as no hyperparameter search was performed for a fair comparison.
Table 6: Quantitative summary - classes ViT-384. The table presents the metrics consistently with those discussed for EfficientNet.
Domain Metric DSEG SLIC QS FS WS
L S G B L S G B L S G B L S G B L S G B
Correctness Random Model $\uparrow$ 84 85 86 85 87 88 87 87 89 89 90 89 88 87 86 86 86 86 87 86
Random Expl. $\uparrow$ 93 94 96 95 97 95 94 96 94 94 99 96 95 95 98 98 94 94 91 95
Single Deletion $\uparrow$ 43 43 44 43 21 22 21 24 20 19 18 17 20 20 22 20 18 17 19 20
Output Completeness Preservation $\uparrow$ 37 35 35 36 27 27 29 28 24 25 25 25 22 20 24 20 23 24 25 25
Deletion $\uparrow$ 82 84 84 83 74 74 73 73 62 61 62 61 63 64 63 66 67 66 67 66
Consistency Noise Preservation $\uparrow$ 32 30 29 32 27 28 29 27 23 21 22 23 22 21 21 22 23 25 24 26
Noise Deletion $\uparrow$ 82 82 82 84 71 73 74 74 67 67 66 65 61 62 61 61 66 65 65 67
Contrastivity Preservation $\uparrow$ 63 61 63 64 56 53 54 54 42 47 46 43 59 58 60 59 46 50 49 47
Deletion $\uparrow$ 69 68 72 69 46 46 46 46 40 40 41 40 61 62 61 61 42 42 43 42
Table 7: Quantitative summary - numbers ViT-384. The table presents the metrics consistently with those discussed for EfficientNet.
Metric DSEG SLIC QS FS WS
L S G B L S G B L S G B L S G B L S G B
Gini $\uparrow$ 0.52 0.52 0.53 0.52 0.48 0.49 0.48 0.47 0.40 0.41 0.42 0.40 0.44 0.42 0.44 0.43 0.43 0.42 0.44 0.42
Incr. Deletion $\downarrow$ 1.00 0.46 0.48 0.51 0.93 0.94 0.91 0.91 1.76 1.72 1.71 1.72 1.64 1.62 1.54 1.59 1.02 1.04 1.03 1.03
Compactness $\downarrow$ 0.20 0.20 0.20 0.19 0.15 0.14 0.15 0.15 0.12 0.12 0.12 0.12 0.14 0.14 0.14 0.14 0.11 0.11 0.11 0.12
Rep. Stability $\downarrow$ .013 .013 .013 .013 .015 .015 .016 .015 .018 .018 .019 .018 .017 .017 .017 .016 .017 .017 .018 .017
Time $\downarrow$ 6.6 6.5 6.6 6.6 3.3 3.2 3.3 3.3 12.2 12.1 12.3 12.3 2.6 2.6 2.5 2.7 3.6 3.5 3.9 3.6
B.3 ConvNext
Table˜8 and Table˜9 present the quantitative results pertaining to the ConvNext model [33], which is the last model that has been thoroughly evaluated in this study. The settings utilized in this evaluation remain consistent with those employed previously.
Table 8: Quantitative summary - classes ConvNext-224. The table presents the metrics consistently with those discussed for EfficientNet.
Domain Metric DSEG SLIC QS FS WS
L S G B L S G B L S G B L S G B L S G B
Correctness Random Model $\uparrow$ 92 93 93 92 81 80 81 81 79 77 79 79 82 82 81 82 83 82 82 82
Random Expl. $\uparrow$ 89 89 90 94 88 87 93 91 94 91 90 93 91 90 90 88 89 86 92 91
Single Deletion $\uparrow$ 44 45 45 44 31 31 29 30 20 20 21 19 27 26 26 27 15 16 16 16
Output Completeness Preservation $\uparrow$ 46 45 46 46 43 43 43 42 27 28 28 28 37 38 37 36 35 35 33 35
Deletion $\uparrow$ 67 66 66 66 64 66 63 63 62 61 55 55 55 55 55 56 53 54 54 54
Consistency Noise Preservation $\uparrow$ 40 47 47 47 39 40 38 41 36 35 38 37 37 38 38 40 40 39 40 38
Noise Deletion $\uparrow$ 68 70 70 69 63 63 64 62 56 56 56 55 57 56 56 58 58 57 59 58
Contrastivity Preservation $\uparrow$ 58 62 61 62 59 58 60 59 50 50 49 46 51 53 52 51 55 56 53 54
Deletion $\uparrow$ 58 58 58 58 47 47 47 47 44 43 44 43 37 37 37 37 46 46 46 47
Table 9: Quantitative summary - numbers ConvNext-224. The table presents the metrics consistently with those discussed for EfficientNet.
Metric DSEG SLIC QS FS WS
L S G B L S G B L S G B L S G B L S G B
Gini $\uparrow$ 0.48 0.49 0.49 0.49 0.51 0.50 0.51 0.51 0.43 0.44 0.45 0.44 0.47 0.47 0.48 0.49 0.48 0.48 0.48 0.48
Incr. Deletion $\downarrow$ 0.87 0.35 0.36 0.37 0.61 0.60 0.60 0.60 1.13 1.14 1.14 1.13 1.10 1.10 1.10 1.10 0.65 0.64 0.65 0.65
Compactness $\downarrow$ 0.23 0.23 0.23 0.23 0.19 0.19 0.19 0.19 0.15 0.15 0.14 0.15 0.16 0.16 0.16 0.16 0.16 0.16 0.16 0.16
Rep. Stability $\downarrow$ .012 .012 .013 .012 .011 .011 .012 .012 .012 .012 .013 .012 .012 .012 .013 .012 .011 .011 .011 .011
Time $\downarrow$ 3.6 3.4 3.5 3.6 1.0 0.9 1.0 1.0 2.8 2.5 2.5 2.9 1.5 1.4 1.5 1.5 0.9 0.9 0.9 1.0
B.4 DSEG compared to EAC
We conducted additional experiments with Explain Any Concept (EAC) [52], performing the same quantitative experiments as for DSEG, but with a reduced dataset comprising 50 images to optimize computational time (Section˜D.1). We began our evaluation by noting that EAC, unlike DSEG-LIME, cannot be applied to arbitrary models, which is a significant drawback of their method and prevents comprehensive comparisons. Thus, we compared our approach against LIME and EAC, in explaining ResNet. The results are listed in Table˜10 and Table˜11.
Table 10: Quantitative summary - classes EAC.This table presents the metrics in line with the previous evaluations, focusing on ResNet performance for DSEG and other segmentation techniques in comparison to EAC.
Domain Metric LIME-DSEG LIME-SLIC LIME-QS LIME-FS LIME-WS EAC
Correctness Random Model $\uparrow$ 45 42 44 40 46 45
Random Expl. $\uparrow$ 49 45 45 46 44 48
Single Deletion $\uparrow$ 10 7 4 7 8 1
Output Completeness Preservation $\uparrow$ 20 21 12 15 19 35
Deletion $\uparrow$ 27 18 20 18 18 38
Consistency Noise Stability $\uparrow$ 20 17 13 12 18 34
Contrastivity Preservation $\uparrow$ 17 13 19 19 21 31
Deletion $\uparrow$ 25 23 24 22 18 35
We observe that EAC quantitatively outperforms DSEG in certain cases. However, the results indicate that DSEG shows marked improvement as the number of samples increases, ultimately achieving comparable computation times. Moreover, it is expected that EAC performs better with ResNet, as it is specifically designed to leverage the model’s internal representations. The main drawback of EAC, however, is its lack of general applicability, as it cannot be used across all model architectures.
Table 11: Quantitative summary - numbers EAC. This table presents the metrics in line with the previous evaluations, focusing on ResNet performance for DSEG and other segmentation techniques in comparison to EAC.
Metric LIME-DSEG LIME-SLIC LIME-QS LIME-FS LIME-WS EAC
Incr. Deletion $\downarrow$ 0.54 0.55 0.90 0.85 0.50 0.01
Compactness $\downarrow$ 0.25 0.16 0.14 0.16 0.13 0.11
Rep. Stability $\downarrow$ .021 .019 .018 .018 .017 .002
Time $\downarrow$ 8.0 2.8 12.6 2.9 3.5 326.7
B.5 DSEG within SLICE
Table˜12 reports the results for the DSEG framework applied to the SLICE [7] approach and conducts a comparative evaluation against Quickshift (QS) in LIME. The experiments were carried out on the minimized dataset consisting of 50 images in total. Each image was processed for 200 steps and repeated 8 times. The primary evaluation metric used was the average rank similarity (p = 0.3), which focused on positive and negative segments and ensured comparability by employing the minimum segment count of both approaches. Furthermore, a sign flip analysis was performed, measuring the number of segments that changed signs during the computation. This analysis considered all segments produced by each technique.
Table 12: DSEG and QS in SLICE. The Average Rank Similarity shows high similarity for both approaches, while the Sign Flip metric indicates that DSEG produces fewer segment sign changes.
Metric DSEG (pos, neg) QS (pos, neg)
Average Rank Similarity (p=0.3) 0.969347 / 0.971652 0.970756 / 0.970384
Sign Flip (mean) 2.608696 31.288043
DSEG and QS demonstrate comparable performance in rank similarity. However, DSEG markedly surpasses QS in the Sign Flip metric. This superior performance is attributable to QS producing a greater number of segments, whereas DSEG prioritizes generating meaningful segments that align more closely with the foundational model. These findings underscore DSEG’s advantage in necessitating less hyperparameter tuning and yielding more robust segmentation compared to alternative methodologies.
B.6 DETR within DSEG
In Table˜13 and Table˜14, we conducted the DETR experiments within LIME. Based on prior results, we assess its performance by contrasting it with SLIC within the LIME framework, also utilizing the same condensed dataset comprising 50 images. Both experiments were configured with identical parameters, and DETR was implemented for basic panoptic segmentation.
Table 13: Quantitative summary - classes DETR. The table showcases metrics for EfficientNetB4, and DETR within DSEG. Results reported pertain solely to integrating DSEG and SLIC within the scope of the LIME frameworks examined.
Domain Metric DSEG SLIC
L S G B L S G B
Correctness Random Model $\uparrow$ 32 32 32 32 30 30 30 30
Random Expl. $\uparrow$ 29 37 38 40 38 45 39 38
Single Deletion $\uparrow$ 36 36 35 36 18 17 21 21
Output Completeness Preservation $\uparrow$ 43 42 42 42 37 35 35 35
Deletion $\uparrow$ 34 34 35 34 21 21 21 21
Consistency Noise Stability $\uparrow$ 40 40 39 39 35 36 36 36
Contrastivity Preservation $\uparrow$ 39 37 36 36 28 28 27 28
Deletion $\uparrow$ 35 34 33 32 23 24 24 24
Table 14: Quantitative summary - numbers DETR. The table showcases the numeric values in the same manner as in Table˜13 but for numeric values.
Metric DSEG SLIC
L S G B L S G B
Incr. Deletion $\downarrow$ 0.64 0.34 0.37 0.25 0.68 0.70 0.75 0.69
Compactness $\downarrow$ 0.34 0.34 0.34 0.34 0.15 0.14 0.15 0.15
Rep. Stability $\downarrow$ .008 .008 .008 .007 .010 .010 .011 .010
Time $\downarrow$ 23.6 22.0 24.4 23.5 22.9 24.5 27.6 25.6
DETR demonstrates superior performance on the dataset compared to the LIME variants utilizing SLIC. Despite its efficacy, the segmentation quality of DETR was generally inferior to that of SAM, as evidenced by less compact explanations. This observation is further supported by the examples in Figure˜5. The visualizations reveal that DETR often segments images in ways that do not align with typical human-recognizable concepts, highlighting a potential limitation in its practical utility for generating explanatory segments. Moreover, DETR does not support the construction of a segmentation hierarchy, lacking the ability to produce finer and coarser segments, which diminishes its flexibility compared to methods such as SAM.
Figure 5: DETR within DSEG. The visualization displays five instances with classes from the ImageNet dataset. Each image includes the prediction by EfficientNetB4 as its headline, the segmentation map of DETR, and the corresponding explanation by DETR within LIME.
B.7 SAM 2 within DSEG
In the main paper, we conducted experiments using SAM 1. In this part, we integrate SAM 2 [48] with the ’hiera_l’ backbone into the DSEG framework, applying a 0.8 stability score threshold. We also report the results for EfficientNetB4 on the dataset comprising 50 images in Table˜15 and Table˜16.
Table 15: Quantitative summary - classes SAM 2. This table presents the metrics in line with those discussed for EfficientNet in the main paper, but displays only the results for SLIC for the sake of simplicity.
Domain Metric DSEG SLIC
L S G B L S G B
Correctness Random Model $\uparrow$ 38 38 38 38 30 30 30 30
Random Expl. $\uparrow$ 40 44 43 44 38 45 39 38
Single Deletion $\uparrow$ 27 27 28 28 18 17 21 21
Output Completeness Preservation $\uparrow$ 39 34 35 34 37 35 35 35
Deletion $\uparrow$ 34 30 30 30 21 21 21 21
Consistency Noise Stability $\uparrow$ 38 38 38 37 35 36 36 36
Contrastivity Preservation $\uparrow$ 31 28 29 30 28 28 27 28
Deletion $\uparrow$ 35 30 31 31 23 24 24 24
As both tables demonstrate, DSEG-LIME consistently outperforms other methods and surpasses DSEG with SAM 1 across most metrics, delivering superior results. It effectively segments images into more meaningful regions, particularly in cases where SAM 1 faced challenges, reinforcing the conclusions of the SAM 2 technical report.
Table 16: Quantitative summary - numbers SAM 2. This table presents the metrics in line with those discussed for EfficientNet in the main paper, but displays only the results for SLIC for the sake of simplicity.
Metric DSEG SLIC
L S G B L S G B
Incr. Deletion $\downarrow$ 0.98 0.36 0.36 0.39 0.68 0.70 0.75 0.69
Compactness $\downarrow$ 0.16 0.16 0.17 0.17 0.15 0.14 0.15 0.15
Rep. Stability $\downarrow$ .011 .010 .011 .010 .010 .010 .011 .010
Time $\downarrow$ 19.1 18.9 18.7 19.3 14.7 14.6 14.8 14.8
However, since the experiments were conducted on different hardware, the computation times vary. Here, we report the time for the SLIC variant of LIME, but similar to previous experiments, the times for other LIME variants with SLIC are expected to be comparable to those of standard LIME. As a result, DSEG with SAM 2 is slightly slower due to the additional segmentation process.
Exemplary explanations. Figure˜6 presents explanations generated by DSEG using both SAM 1 and SAM 2, highlighting cases where the newer version of SAM enables DSEG to produce more meaningful and interpretable explanations. Each image includes explanations for the predicted class from EfficientNetB4. While SAM 2 shows improved segmentation in these examples, similar results can be obtained with SAM 1 by appropriately adjusting the hyperparameters for automatic mask generation.
Figure 6: Comparison DSEG with SAM 1 and SAM 2. Exemplary images with explanations generated by SAM 1 and SAM 2 within DSEG, illustrating how the updated SAM improves segment utilization for DSEG.
B.8 EfficientNetB4 with depth of two
In Table˜17 and Table˜18, we present the quantitative comparison between DSEG-LIME ($d=2$) using EfficientNetB4 and SLIC, as reported in the main paper. The hyperparameter settings were consistent across the evaluations, except for compactness. We established a minimum threshold of 0.05 for values to mitigate the impact of poor segmentation performance, which often resulted in too small segments. Additional segments were utilized to meet this criterion for scenarios with suboptimal segmentation. However, this compactness constraint was not applied to DSEG with depth two since its hierarchical approach naturally yields smaller and more detailed explanations, evident in Table˜18. The hierarchical segmentation of $d=2$ slightly impacts stability, yet the method continues to generate meaningful explanations, as indicated by other metrics. Although our method demonstrated robust performance, it required additional time because the feature attribution process was conducted twice.
Table 17: Quantitative summary - classes depth two. The table showcases metrics for EfficientNetB4, specifically at a finer concept granularity; the hierarchical segmentation tree has $d=2$. Results reported pertain solely to integrating DSEG and SLIC within the scope of the LIME frameworks examined.
Domain Metric DSEG SLIC
L S G B L S G B
Correctness Random Model $\uparrow$ 62 68 69 69 68 67 68 68
Random Expl. $\uparrow$ 87 87 87 91 81 79 80 84
Single Deletion $\uparrow$ 38 40 45 43 36 34 33 34
Output Completeness Preservation $\uparrow$ 63 70 68 68 74 74 75 74
Deletion $\uparrow$ 45 49 50 52 39 39 39 40
Consistency Noise Preservation $\uparrow$ 65 63 63 64 72 72 72 72
Noise Deletion $\uparrow$ 54 52 53 52 40 41 40 40
Contrastivity Preservation $\uparrow$ 51 53 52 53 54 54 55 55
Deletion $\uparrow$ 50 46 49 48 49 50 48 50
Table 18: Quantitative summary - numbers depth two. The table showcases the numeric values in the same manner as in Table˜17 but for numeric values.
Metric DSEG SLIC
L S G B L S G B
Gini $\uparrow$ 0.42 0.43 0.42 0.49 0.49 0.50 0.50 0.50
Incr. Deletion $\downarrow$ 1.29 0.85 0.80 1.10 0.81 0.82 0.81 0.82
Compactness $\downarrow$ 0.16 0.17 0.17 0.16 0.15 0.15 0.15 0.15
Rep. Stability $\downarrow$ .014 .015 .016 .015 .011 .011 .012 .012
Time $\downarrow$ 47.6 52.5 53.4 52.8 22.9 24.5 27.6 25.6
Exemplary explanations. In Figure˜7, an instance is explained with DSEG and $d=2$, showing a black-and-white image of a projectile. Here, we see the corresponding explanation for each stage, starting with the first iteration, with the corresponding segmentation map. In the second iteration, we see the segment representing the projectile split into its finer segments - the children nodes of the parent node - with the corresponding explanation below.
Figure 7: 2nd iteration of DSEG-LIME. Visualizing DSEG’s explanations of a projectile. It includes the first iteration’s explanation along with its corresponding segmentation map. Additionally, similar details are provided for the second iteration procedure, highlighting the upper part of a projectile as an explanation.
Case study. We examine the case presented in Figure˜3, where DSEG initially segments the image into various layers with overlapping features, establishing a segmentation hierarchy through composition. In the first iteration, LIME focuses solely on the segments just beneath the root node - the parent segments that cannot be merged into broader concepts. From this segmentation map, LIME determines the feature importance scores, identifying the airplane as the most crucial element in the image. In the subsequent iteration, illustrated in Figure˜8, DSEG generates an additional segmentation map that further divides the airplane into finer components for detailed analysis. The explanation in this phase emphasizes the airplane’s body, suggesting that this concept of the ’Airliner’ is most significant.
Figure 8: Airliner explanation with depth two. The same example as in Figure˜3 but with segmentation hierarchy of two for the explanation. This example includes the children nodes of the most significant parent node in the segmentation map for feature importance calculation.
Appendix C Further experiments with DSEG
C.1 Ablation study
Automatic mask generation. Ablating DSEG using only the automatic masks produced by foundation models presents notable challenges, particularly when applying methods like LIME. These automatically generated masks often result in many segments, some of which may overlap or leave parts of the image unsegmented. Such inconsistencies make a direct integration with LIME impractical. Consequently, our ablation in this setting is limited to comparing different foundation models for mask generation. We conducted experiments with SAM, SAM2, and DETR to evaluate their respective impact on DSEG’s performance.
Small cluster removal. For the ablation study, we examine how the number of segments evolves across different stages of the segmentation process as we vary the threshold for removing segments smaller than the hyperparameter $\theta$ (with values [100, 300, 500, 1000, 2000]). Additionally, we assess the behavior of empty spaces within the segmented regions across all images in the dataset. The analysis focuses on three key points: the number of segments immediately after the initial automated segmentation, after hierarchical sorting, and after the removal of undersized segments, following the complete DSEG approach. The empty space is evaluated before it is filled with adjacent segments. A comprehensive overview of the metrics for these steps is presented in Figure˜9.
Figure 9: Ablation study. Here we present the interquartile range (IQR) of segmentation counts at different stages of the DSEG process (before hierarchy, after hierarchy, and final segmentation) and the proportion of empty space across various threshold values for segment size removal (denoted by $\theta$).
Higher thresholds lead to fewer segments being retained. This trend is visible in the segmentation counts before the hierarchy, after the hierarchy, and in the final segmentation. For instance, at $\theta=100$, a higher number of segments is preserved, whereas at $\theta=2000$, the segmentation count drops significantly due to the removal of smaller segments. Additionally, the proportion of empty space consistently increases with larger $\theta$ values. This occurs because as more small segments are removed, more unassigned or empty regions appear before being filled by adjacent segments. The increase in empty space proportion is most pronounced at higher thresholds, such as $\theta=1000$ and $\theta=2000$. In summary, the analysis highlights the expected trade-off between preserving smaller segments and controlling the amount of empty space. Lower thresholds result in more granular segmentation, while higher thresholds reduce the segmentation complexity at the expense of increased empty regions. Based on this trade-off, a threshold of $\theta=500$ was selected for the experiments in this paper, as it strikes a balance between retaining meaningful segmentation detail and minimizing empty space.
Hierarchical ordering. The effect of hierarchical ordering is already addressed through both qualitative and quantitative analyses across different depths of the explanation process. This hierarchical structure ensures no overlapping segments occur—a necessary condition for reliably computing feature attributions. As a result, an additional technical ablation isolating hierarchical ordering is not required.
Empty space removal. The presence of unsegmented regions within an image is a well-known issue in the segmentation domain. One potential solution for adapting LIME to this scenario is to assign these unsegmented areas to a dedicated segment—e.g., a “segment zero.” However, this approach introduces the challenge that such unsegmented areas are often non-contiguous and scattered across the image. From a user perspective, this fragmented behavior is undesirable and could negatively affect interpretability. To address this, we opted for a simple yet practical solution: reassigning unsegmented pixels to the nearest segmented region using a nearest-neighbor approach. While effective for our purposes, this strategy opens avenues for future work, where more sophisticated treatments of unsegmented areas could be explored.
C.2 Medical scenario with depth of three
Figure 10: Medical use case. Demonstration of DSEG on a medical use case using CT scan imagery. From left to right: the original CT image, the raw segmentation mask, and hierarchical segmentations generated by DSEG at depths 1, 2, and 3. Each subsequent depth introduces finer segmentation details important for explanation. This showcases the general motivation for hierarchical explanations and the applicability of DSEG in medical imaging contexts.
We demonstrate the application of our framework on a medical use case, processing CT imagery of the human body, as shown in Figure˜10. In this example, we present the original image alongside the raw segmentation mask and DSEG’s hierarchical compositions at depths 1 to 3, where each level introduces progressively finer segments relevant for explanation.
For this demonstration, we utilize the basic SAM variant. However, improved segmentation quality could be achieved by using fine-tuned models specifically adapted to the medical domain. It is important to note that we do not provide a detailed explanation of a specific instance here; rather, we aim to illustrate the general motivation behind DSEG. Thus, the presented hierarchy serves demonstration purposes but highlights the feasibility and applicability of our framework when paired with appropriate segmentation models for medical data.
Furthermore, we show that the segmentation hierarchy can extend beyond depth 2, as previously evaluated in Section˜B.8. As a final remark, if a user requests an explanation deeper than the maximum available hierarchy (determined by the most activated nodes), the explanation terminates due to the inherent limitation in segmentation depth.
C.3 Explaining wrong classification
Here, we explore how DSEG can aid in explaining a model’s misclassification. Unlike the previous analysis in Section˜5.2.1, where metrics were assessed under simulations involving a model with randomized weights (Random Model) or random predictions (Random Expl.), this case focuses on a real misclassification by EfficientNetB4, free from external manipulation. This allows for a more genuine examination of DSEG’s ability to explain incorrect classifications under normal operating conditions.
Figure 11: Misclassification example. The image depicts a hybrid of a horse and zebra that EfficientNetB4 classifies as a zebra with $p=0.17$. DSEG-LIME, with a depth of one, highlights the entire animal, offering a broad explanation. Meanwhile, DSEG at depth two pinpoints specific zebra-like patterns that influence the model’s prediction. This suggests that the model is fixating on particular visual features associated with zebras, explaining its erroneous classification.
Figure˜11 shows an image of a hybrid between a horse (sorrel) and a zebra, where EfficientNetB4 can recognize both animals but does not contain the hybrid class. We explore why EfficientNetB4 assigns the highest probability to the zebra class rather than the sorrel. Although this is not strictly a misclassification, it simulates a similar situation and provides insight into why the model favors the zebra label over the sorrel. This analysis helps us understand the model’s decision-making process in cases where it prioritizes specific features associated with one class over another.
C.4 Stability of explanations
Figure 12: Segmentation stability. Illustrating a comparison between DSEG and other segmentation techniques applied in LIME, all utilizing an identical number of samples. DSEG exhibits greater stability compared to other segmentation techniques. Notably, the DSEG explanation distinctly highlights the segment representing a gorilla as the most definitive.
The stability of imagery explanations using LIME can be linked to the quality of feature segments, as illustrated in Figure˜12. This figure presents the segmentation maps generated by various techniques alongside their explanations and coefficient distributions, displayed through an IQR plot over eight runs. Notably, the DSEG technique divides the image into meaningful segments; the gorilla segment, as predicted by the EfficientNetB4 model, is distinctly visible and sharply defined. In contrast, other techniques also identify the gorilla, but less clearly, showing significant variance in their coefficient distributions. Watershed, while more stable than others, achieves this through overly broad segmentation, creating many large and a few small segments. These findings align with our quantitative evaluation and the described experimental setup.
C.5 Zero-shot classification explanation
In this section, we demonstrate the versatility of DSEG-LIME by applying it to a different dataset and classification task. Specifically, we replicate the zero-shot classification approach described in [43] using CLIP [44] for the animal super-category. Since DSEG-LIME maintains model-agnostic properties, it remains applicable to zero- and few-shot classification models without modification.
Figure˜13 presents an illustrative example from the dataset, where the task is to classify an image into the animal category. The predicted and ground-truth class for the image is ’Land mammal’. As shown by DSEG-LIME’s explanation, the model’s decision is primarily influenced by the presence of a deer in the foreground and a mountain in the background, which contribute to the overall classification.
Figure 13: DSEG-LIME explanation for CLIP. This figure illustrates an image processed by CLIP for zero-shot classification into animal categories. The model correctly predicted the class as "Land mammal." DSEG-LIME highlights the two most important features influencing the classification, with the presence of the deer being the most significant.
C.6 Exemplary limitation of DSEG
The example in Figure˜14 shows a complex case of a hermit crab in front of sand, which is hardly detectable. Here, SAM fails to segment the image into meaningful segments, a known issue in the community [26]. In contrast, SLIC can generate segments; thus, LIME can produce an explanation that does not show a complete image. However, to mitigate this issue, one could use a segmentation model that has better performance in such scenes, such as [10].
Figure 14: DSEG fails. Demonstrating a scenario where DSEG fails to generate meaningful features for explanations (the the whole image is one segment, in contrast to SLIC. The image shows a crab, which the model classifies as a ’hermit crab’ (p = 0.17), highlighting the effectiveness of SLIC in this context compared to the limitations of DSEG.
C.7 Feature attribution maps
In addition to visualizing the $n$ most essential segments for an explanation, feature attribution maps also help the explainee (the person receiving the explanation [37]) to get an idea of which other segments are important for interpreting the result. In these maps, the segments represent the corresponding coefficient of the surrogate model learned within LIME for the specific case. Figure˜15 represents all feature maps of EfficientNetB4 with the reported settings, accompanied by the original image with the most probable class. Blue segments are positively associated with the class to be explained, and red segments are negatively associated.
Figure 15: DSEG attribution maps. Representation of the feature weights of all 100 images, with blue segments indicating positively important and red segments negatively important features in relation to the classified label.
Appendix D Dataset and user study
D.1 Dataset
Image selection. As mentioned in Section˜5.1, we selected various classes of images from the ImageNet [13] and COCO [31] dataset. Additionally, we created artificial images using the text-to-image model DALL-E [46] to challenge the XAI techniques when facing multiple objects. The dataset for the evaluation comprised 97 real images and three synthetic images. For the synthetic instances, the prompts ’realistic airplane at the airport’ (Airplane), ’realistic person running in the park’ (Miniskirt), and ’realistic person in the kitchen in front of a dishwasher’ (Dishwasher) were used.
The object types listed in Table˜19 represent the primary labels of the images used in the dataset. Each image is unique, ensuring no duplication and maximizing the diversity of animals and objects covered. The bolded types denote those that were randomly selected for qualitative evaluation. The cursive types signify the reduced dataset comprising fifty instances, implemented to conserve computational time for supplementary evaluations. These types provide a balanced representation of the dataset and were chosen to ensure broad coverage across different categories. This selection strategy helps to avoid bias and supports a comprehensive evaluation.
Table 19: Object families and types. This table categorizes the images in the dataset according to their object families. The bold text denotes the classes selected for the user study, while the italicized text represents the minimized dataset; both were chosen at random to ensure objectivity variety.
Object Family Type
Animals Ice_bear, Gorilla, Chihuahua, Husky, Horse, Irish_terrier, Macaw,
American_lobster, Kerryblue_terrier, Zebra, House_finch, American_egret,
Little_blue_heron, Tabby, Black_bear, Egyptian_cat, Tusker, Quail, Affenpinscher
Leatherback_turtle, Footed_ferret, Howler_monkey, Blenheim_spaniel, Otter
Silky_terrier, Cocker_spaniel, Hare, Siberian_husky, Harvestman, Sea_snake
Cock, Scottish_deerhound, Tiger_cat, Hyena
Objects Street_sign, Park_bench, CD_player, Banana, Projectile, Ski,
Catamaran, Paper_towel, Violin, Miniskirt, Basketball, Tennis_racket,
Airplane, Dishwasher, Scuba_diver, Pier, Mountain_tent, Totem_pole,
Bullet_train, Lakeside, Desk, Castle, Running_shoes, Snorkel, Digital_Watch,
Church, Refrigerator, Meat_loaf, Dome, Forklift, Teddy, Mosque, Shower_curtain
Four_poster, Photocopier, Stone_wall, Crib, Bow_tie, Measuring_cup, Unicycle
Cowboy_hat, Dutch_oven, Go-kart, Necklace, Bearskin, Sleeping_bag, Trombone
Microphone, Sandal, Fireboat, Carousel, Drum, Shoji, Solar_dish, Stove, Cup
Panpipe, Custard_apple, Gondola, Minivan, Bagel, Lighter, Pot, Carton, Ear
Volleyball, Plow, Mailbox, Bucket, Chime, Toaster
D.2 User study
We conducted our research and user study using MTurk, intentionally selecting participants without specialized knowledge to ensure the classes represented everyday situations. Each participant received compensation of \$4.50 per survey, plus an additional \$2.08 handling fee charged by MTurk and \$1.24 tax. The survey, designed to assess a series of pictures, takes approximately 10 to 15 minutes to complete. The sequence in which the explanations are presented to the participants was randomized to minimize bias. In our study conducted via MTurk, 59 individuals participated, along with an additional 28 people located near our research group who participated at no cost.
Explanations. In Figures˜18 and 18 we show all 20 images from the dataset used for the qualitative evaluation. Each image is accompanied by the prediction of EfficientNetB4 and the explanations within the vanilla LIME framework with all four segmentation approaches and the DSEG variant. The segments shown in the image indicate the positive features of the explanation.
Figure 16: Images 1–10.
Figure 17: Images 11–20.
Figure 18: User study data. Examples from the evaluation datasets showing the LIME explanations alongside the original images and their corresponding predictions.
Figure 19: Exemplary question. The ’airplane’ example is shown in the original image with its five explanations. Below the images, participants can rate the quality of the explanations accordingly.
Instruction. Participants were tasked with the following question for each instance: ’Please arrange the provided images that best explain the concept [model’s prediction], ranking them from 1 (least effective) to 5 (most effective).’ Each instance was accompanied by DSEG, SLIC, Watershed, Quickshift, Felzenszwalb, and Watershed within the vanilla LIME framework and the hyperparameters discussed in the experimental setup. These are also the resulting explanations used in the quantitative evaluation of EfficientNetB4. Figure˜19 shows an exemplary question of an instance of the user study.
Results. We show the cumulative maximum ratings in Figure˜22 and in Figure˜22 the median (in black), the interquartile range (1.5), and the mean (in red) for each segmentation technique. DSEG stands out in the absolute ratings, significantly exceeding the others. Similarly, in Figure˜22, DSEG achieves the highest rating, indicating its superior performance relative to other explanations. Therefore, while DSEG is most frequently rated as the best, it consistently ranks high even when it is not the leading explanation, as the IQR of DSEG shows. Aligned with the quantitative results in Section˜5.2, the Quickshift algorithm performs the worst.
Figure 20: Best rated explanation. Accumulated number of best-selected explanations within the user study. DSEG was selected as the favourite, followed by Felzenszwalb and Watershed.
Figure 21: IQR of explanation’s ratings. The IQR plot of the user study ratings is detailed, with the black line indicating the median and the red line representing the mean. This plot shows that DSEG received the highest ratings, while Watershed exhibited the broadest ratings distribution.
Figure 22: User study results. The user study ratings are visualized in two distinct figures, each employing a different form of data representation. In both visualizations, DSEG consistently outperforms the other techniques.
Table˜20 presents the statistical significance of the user study. Specifically, it lists the t-statistics and p-values for comparisons between DSEG (the baseline method) and other segmentation methods, namely SLIC, QS, FS, and WS. The t-statistics indicate the magnitude of difference between DSEG and each different method, with higher values representing greater differences. The corresponding p-values demonstrate the probability that these observed differences are due to random chance, with lower values indicating stronger statistical significance.
Table 20: User study statistical results. This table summarizes the statistical significance of user study results for each segmentation approach. The t-statistics and p-values indicate the comparison between DSEG and other methods. Extremely low p-values suggest strong statistical significance.
Metric DSEG SLIC QS FS WS
t-statistics $\uparrow$ 20.01 49.39 20.89 33.15
p-values $\downarrow$ 8.0e-143 << 2.2e-308 1.2e-86 3.3e-187
In this context, the null hypothesis (H0) posits no significant difference in participant preferences between the performance of DSEG and other segmentation methods within the LIME framework. The alternative hypothesis (HA) asserts that DSEG performs significantly better than the different segmentation methods on the dataset when evaluated using five explanations. Given the extremely low p-values (e.g., 8.0e-143 for SLIC and < 2.2e-308 for QS), we can reject the null hypothesis (H0) with high confidence. The significance level of 99.9% ($\alpha$ = 0.001) further supports this conclusion, as all p-values fall well below this threshold. These results indicate that the observed differences are highly unlikely to have occurred by chance and are statistically significant.